OpenStudy (anonymous):

find the solution to the equation: -2(x+7)+2x=5x+3

jimthompson5910 (jim_thompson5910):

-2(x+7)+2x=5x+3 -2x-14+2x=5x+3 -14=5x+3 0=5x+3+14 0-5x=3+14 -5x=3+14 -5x=17 x=(17)/(-5) x=-17/5 So the answer is \[\large x=-\frac{17}{5}\]